# Are We Truly Forgetting? A Critical Re-examination of Machine Unlearning Evaluation Protocols

### v1

@@ Abstract
Machine unlearning is a process to remove specific data points from a trained model while maintaining the performance on retain data, addressing privacy or legal requirements. Despite its importance, existing unlearning evaluations tend to focus on logit-based metrics (i.e., accuracy) under small-scale scenarios. We observe that this could lead to a false sense of security in unlearning approaches under real-world scenarios. In this paper, we conduct a new comprehensive evaluation that employs representation-based evaluations of the unlearned model under large-scale scenarios to verify whether the unlearning approaches genuinely eliminate the targeted forget data from the model’s representation perspective. Our analysis reveals that current state-of-the-art unlearning approaches either completely degrade the representational quality of the unlearned model or merely modify the classifier (i.e., the last layer), thereby achieving superior logit-based evaluation metrics while maintaining significant representational similarity to the original model. Furthermore, we introduce a novel unlearning evaluation setup from a transfer learning perspective, in which the forget set classes exhibit semantic similarity to downstream task classes, necessitating that feature representations diverge significantly from those of the original model. Our comprehensive benchmark not only addresses a critical gap between theoretical machine unlearning and practical scenarios, but also establishes a foundation to inspire future research directions in developing genuinely effective unlearning methodologies.
@@ 1 Introduction
Neural networks have driven transformative advancements across a multitude of domains, such as image classification [10, 17, 25]. A key contributor to this progress has been the substantial enhancement of the model capability, facilitated by the training of sophisticated models on vast datasets [1]. However, these successes have also raised significant privacy concerns in these datasets, particularly related to the unchecked use of sensitive data [4]. As privacy rights, including the ‘right to be forgotten’ (e.g., the GDPR legislation [3, 27]), have gained increasing recognition, the need for effective solutions has become urgent. As a result, this has led to the emergence of machine unlearning as a critical area of research [28].
The primary objective of machine unlearning is to effectively eliminate the influence of specific data from a pre-trained model while preserving the knowledge derived from other data [28]. The most straightforward approach, known as exact unlearning, involves retraining the model from scratch using a refined dataset [26] (hereafter referred to as the “retrained model”). While this approach is regarded as the gold standard because it guarantees complete elimination of any influence from data designated for removal, it incurs significant computational expenses due to the necessity of retraining models on refined datasets whenever unlearning is required. To overcome these limitations, several approximate unlearning algorithms have been proposed, offering more efficient alternatives while striving to maintain the goal of machine unlearning [16]. In image classification tasks, approximate unlearning algorithms are typically evaluated based on their ability to produce models that achieve classification accuracy similar to that of the exact unlearning on the given dataset (e.g., retain and forget sets) [28], as illustrated on the left side of Fig. 1. Following this evaluation method, several methods have induced unlearning through techniques such as using random labeling [14], gradient ascent [32] and metric learning [2], while mitigating the over-forgetting by incorporating additional regularization terms [23, 12, 6, 2].
@@
[…]
@@
Figure 2: Performance comparison between logit-based (left) and representation-based evaluation (right) reveals contrasting findings. In the left figure, logit-based metrics suggest that the unlearned model ($\theta_{u}$) performs similarly to the retrained model ($\theta_{r}$), implying seemingly successful unlearning. However, the right figure, which presents our proposed representation-based evaluation, shows a different trend. Negative values in the right figure indicate that the unlearned model remains closer to the original model ($\theta_{o}$) than to the retrained model ($\theta_{r}$), revealing that none of the unlearning approaches achieve greater similarity to $\theta_{r}$ than to $\theta_{o}$.
@@
In this paper, we propose a comprehensive new evaluation framework that addresses the limitations of existing small-scale, logit-based evaluations by incorporating assessments of unlearned models ($\theta_{u}$) from the perspective of feature representations in large-scale settings. As illustrated in Fig. 1(b), our framework evaluates the representations of unlearned models using Centered Kernel Alignment (CKA) [20] to compare the feature representations of the unlearned models with those of a model trained only on the retain set ($\theta_{r}$). Additionally, we apply $k$-Nearest Neighbors ($k$-NN) analysis on the feature space across various downstream datasets to assess the quality of feature representations. Through these evaluation methods, we re-evaluate multiple state-of-the-art unlearning algorithms in the large-scale unlearning scenario. Fig. 2 illustrates key findings from experiments on unlearning 100 random classes in ImageNet-1k using ResNet-50: while logit-based evaluation (e.g., classification accuracy) demonstrates successful unlearning by showing a decreasing gap between the unlearned models ($\theta_{u}$) and the retrained model ($\theta_{r}$) (Fig. 2, left), a representation-based evaluation using CKA reveals a critical limitation (Fig. 2, right). In an ideal scenario, where unlearning is truly effective, the unlearned model should be indistinguishable from the retrained model, with the ideal value being 10. However, as shown in Fig. 2 (right), all unlearned models exhibit greater similarity to the original model than to the retrained model. This highlights a fundamental issue: existing unlearning methods fail to sufficiently erase learned information at the representation level. This reinforces the need for representation-based evaluation to rigorously assess whether unlearning truly eliminates the influence of the targeted data. Furthermore, random class forgetting has been widely adopted as a standard evaluation scenario in unlearning studies [6, 12, 2, 23]. However, we highlight a critical limitation of this approach: the forgetting process may be inadequately assessed, as similar representational knowledge can persist across multiple alternative classes, making it difficult to verify whether the targeted data has been truly unlearned. To address this issue, we propose a novel unlearning evaluation paradigm, Top Class-wise Forgetting, in which the classes selected for unlearning are semantically similar to those in downstream tasks. This setup requires a substantial divergence of feature representations from the original model, offering a more rigorous assessment of unlearning algorithm’s effectiveness. We then evaluate existing unlearning methods using both logit-based and representation-based metrics across diverse downstream datasets. In summary, the core contributions are summarized as follows:
  • We explore new evaluation metrics for unlearning from a representational perspective under large-scale settings, demonstrating that existing unlearning models exhibit limited effectiveness or fail to genuinely eliminate targeted data.
  • We propose a new evaluation setting, Top Class-wise Forgetting, which overcomes the limitations of random class forgetting and enables a more comprehensive assessment of unlearning through diverse downstream tasks from a transfer learning perspective.
  • Our experimental results reveal that state-of-the-art unlearning algorithms frequently fail to perform effectively under our evaluation framework, particularly when assessed from a feature representation perspective.
@@ 2 Related Works
Machine unlearning in image classification. Early research on machine unlearning in neural networks has primarily focused on image classification, leading to the development of various unlearning algorithms with distinct strategies [28]. A widely used approach, Gradient Ascent (GA), maximizes the loss on the forget set to erase its influence, making it one of the most fundamental yet widely adopted techniques in recent studies [13, 31, 15, 37, 6]. Similarly, Fine-tuning achieves unlearning by inducing catastrophic forgetting on forget set, updating the model solely on the retain set [13, 36]. More sophisticated unlearning techniques have been developed: SCRUB [23] employs a knowledge distillation framework to selectively forget knowledge from the forget set while retaining other knowledge. L2UL [6] enhances unlearning by incorporating two regularization terms, mitigating forgetting at both the feature and weight levels when combined with GA. SalUn [12] modifies the most influential parameters through weight saliency, enabling selective unlearning. SCAR [2] leverages metric learning alongside a distillation-based strategy, aligning forget set representations to the nearest incorrect class using the Mahalanobis distance.
@@
Evaluation and analysis of representations. Transfer learning has become a foundational technique in machine learning, facilitating the adaptation of pre-trained models to diverse downstream tasks [29]. Beyond its practical applications, it is widely used to assess the quality of learned representations. For instance, it has been employed to compare the learned representations of different model architectures  [21] and algorithms across various domains, such as self-supervised learning [8, 18] and continual learning [5]. This evaluation is typically performed using metrics such as $k$-Nearest Neighbors ($k$-NN) and linear probing, which measure how well the learned representations generalize in various downstream tasks. Centered Kernel Alignment (CKA) [20] has been widely used to compare feature representations across models, providing insights into their structural similarities and differences. Furthermore, t-SNE [33] has been extensively utilized to visualize feature representations, offering an intuitive means to compare and analyze how different models organize learned representations. Our paper makes several contributions that distinguish it from previous machine unlearning studies. First, we highlight the issues with the conventional evaluation methods and setups commonly used in machine unlearning studies, such as logit-based evaluation on small-scale datasets. Second, to address these issues, we propose various forms of representation-based evaluation and introduce a more practical unlearning scenario, namely Top Class-wise Forgetting, supported by a range of experiments. Based on experimental findings, we argue that achieving meaningful progress in unlearning research requires a comprehensive re-evaluation of existing evaluation methods and setups.
3 Analysis Setup and Experimental Results
@@
In this section, we present a comprehensive analysis of machine unlearning methods, focusing on both logit-based and representation-based evaluations in the large-scale setup. We review the classical unlearning evaluation protocols and introduce a new evaluation metric and setup in Sec. 3.1. Then, we introduce the motivation of our approach by highlighting limitations in current evaluation frameworks in Sec. 3.2, introduce our proposed evaluation metric in Sec. 3.3, present experimental findings under large-scale settings from a representation perspective in Sec. 3.4, and show results under top class-wise forgetting scenarios from a transfer learning perspective in Sec. 3.5.
3.1 Analysis Setup
@@
Problem setup and objective of machine unlearning. We define the entire dataset as $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, where $x_{i}$ represents the input image, $y_{i}$ denotes its corresponding label, and $N$ is the total number of samples in the dataset. The subset of data targeted for unlearning, referred to as the forget set, is denoted by $\mathcal{D}_{f}\subset\mathcal{D}$, and consists of samples from specific classes that the model is intended to forget. retain set, denoted by $\mathcal{D}_{r}=\mathcal{D}\setminus\mathcal{D}_{f}$, includes all samples in $\mathcal{D}$ that are not part of the forget set. Additionally, we introduce the test forget set, denoted as $\mathcal{D}^{\text{te}}_{f}$, which consists of data points from the same classes as $\mathcal{D}_{f}$. Similarly, the test retain set, denoted as $\mathcal{D}^{\text{te}}_{r}$, contains data points from the same classes as $\mathcal{D}_{r}$. We assume that an original model $\theta$ is trained on the entire dataset $\mathcal{D}$ and is then tasked with unlearning knowledge from $\mathcal{D}_{f}$. Note that the goal of machine unlearning is to produce an unlearned model $\theta_{u}$ that retains the knowledge of $\mathcal{D}_{r}$ while achieving the unlearning of $\mathcal{D}_{f}$. Additionally, we consider a retrained model $\theta_{r}$, which is trained on $\mathcal{D}_{r}$ from scratch, as an oracle for comparison in the unlearning process.
@@
Logit-based evaluation metrics. As a classical evaluation protocol, unlearning models are evaluated on the following metrics: Forget Accuracy (FA): Measures the accuracy of the unlearned model $\theta_{u}$ on $\mathcal{D}_{f}$. Retain Accuracy (RA): Measures the accuracy of the unlearned model $\theta_{u}$ on $\mathcal{D}_{r}$. Test Forget Accuracy (TFA): Classification accuracy on the test forget set $\mathcal{D}^{\text{te}}_{f}$. Similar to FA, this evaluates the model’s ability to forget removed classes beyond the training data. Test Retain Accuracy (TRA): Classification accuracy on the test retain set $\mathcal{D}^{\text{te}}_{r}$. It indicates how effectively the model’s knowledge generalizes to unseen data of the remaining classes. Following [39], we introduce the Average GAP of Logit-based Metrics (AGL), which quantifies the discrepancy between the given model (e.g. the unlearned model $\theta_{u}$) and the golden standard $\theta_{r}$. This metric aggregates the logit-based evaluations:
$AGL=$ $\Bigl{(}1-\mathrm{G}(\theta_{u},\mathcal{D}_{f})\Bigr{)}\times\Bigl{(}1-\mathrm{G}(\theta_{u},\mathcal{D}_{r})\Bigr{)}$ (1)
×$\times\Bigl{(}1-\mathrm{G}(\theta_{u},\mathcal{D}^{\text{te}}_{f})\Bigr{)}\times\Bigl{(}1-\mathrm{G}(\theta_{u},\mathcal{D}^{\text{te}}_{r})\Bigr{)},$
@@
where $\mathcal{D}_{f}$, $\mathcal{D}_{r}$, $\mathcal{D}^{\text{te}}_{f}$, and $\mathcal{D}^{\text{te}}_{r}$ represent the source training datasets and their corresponding test sets, and $\mathrm{G}$ computes the classification accuracy differences,
$\mathrm{G}(\theta_{u},\mathcal{D}_{\text{}})=\bigl{|}\,\text{A}(\theta_{u},\mathcal{D}_{\text{}})-\text{A}(\theta_{r},\mathcal{D}_{\text{}})\bigr{|},$
where $\mathrm{A}$ represents the classification accuracy on $\mathcal{D}$.
@@
[…]
Representation-based evaluation metrics. To tackle limitations of the logit-based evaluation, we propose using representation-based evaluation. Unlike logit-based evaluation, this approach extracts features from the encoder (e.g., after average pooling in ResNet) and measures the similarity and quality of the representations in unlearned models compared to those in the retrained model. We use the following two metrics for representation-based metrics: First, we use Centered Kernel Alignment (CKA) [20] to measure the similarity between the feature representations of $\theta_{u}$ and $\theta_{r}$ as well as $\theta_{u}$ and $\theta_{o}$. Second, we measure $k$-Nearest Neighbors Accuracy ($k$-NN) [9] to assess the representational quality of $\theta_{u}$. We employ a $k$-NN classifier, which takes features of the test data from a model and performs classification based on the neighborhood features. Additionally, we also employ t-SNE [33] to visualize feature distributions. More details on these methods are provided in Fig. A. 2 in the Appendix.
@@
Unlearning baselines. We evaluate several representative machine unlearning algorithms in classification, which can be categorized into two groups: 1) Unlearning without $\mathcal{D}_{r}$, including Gradient Ascent (GA)[32], Random Labeling (RL)[14], Pseudo Labeling (PL)[7], and SalUn[12]; and 2) unlearning with $\mathcal{D}_{r}$, including DUCK [11], Contrastive Unlearning (CU) [38], SCRUB [23], and SCAR [2]. Note that SalUn and SCAR are known as current state-of-the-art methods in each category. A more detailed explanation of each baseline can be found in Sec. A in the Appendix.
Unlearning scenarios. In our experiments and analysis, we consider two unlearning scenarios. The first is Random-$N$ Class-wise Forgetting, where we randomly select $N$ classes to remove. This scenario is a standard approach to evaluate unlearning algorithms. To provide a more comprehensive assessment, we also introduce a second scenario, Top-$N$ Class-wise Forgetting, in which we specifically target classes that exhibit semantic similarity to the downstream task classes. The selection of $N$ classes is based on their relevance to the downstream task.

### v2

@@ Abstract
Machine unlearning is a process to remove specific data points from a trained model while maintaining the performance on retain data, addressing privacy or legal requirements. Despite its importance, existing unlearning evaluations tend to focus on logit-based metrics (i.e., accuracy) under small-scale scenarios. We observe that this could lead to a false sense of security in unlearning approaches under real-world scenarios. In this paper, we conduct a new comprehensive evaluation that employs representation-based evaluations of the unlearned model under large-scale scenarios to verify whether the unlearning approaches genuinely eliminate the targeted forget data from the model’s representation perspective. Our analysis reveals that current state-of-the-art unlearning approaches either completely degrade the representational quality of the unlearned model or merely modify the classifier (i.e., the last layer), thereby achieving superior logit-based evaluation metrics while maintaining significant representational similarity to the original model. Furthermore, we introduce a rigorous unlearning evaluation setup, in which the forgetting classes exhibit semantic similarity to downstream task classes, necessitating that feature representations diverge significantly from those of the original model, thus enabling a more rigorous evaluation from a representation perspective. We hope our benchmark serves as a standardized protocol for evaluating unlearning algorithms under realistic conditions.
[…]
@@ 1 Introduction
Neural networks have driven transformative advancements across a multitude of domains, such as image classification [10, 18, 25]. A key contributor to this progress has been the substantial enhancement of the model capability, facilitated by the training of sophisticated models on vast datasets [1]. However, these successes have also raised significant privacy concerns about these datasets, particularly related to the unchecked use of sensitive data [4]. As privacy rights, including the ‘right to be forgotten’ (e.g., the GDPR legislation [3, 27]), have gained increasing recognition, the need for effective solutions has become urgent. As a result, this has led to the emergence of machine unlearning as a critical area of research [28].
The primary objective of machine unlearning is to effectively eliminate the influence of specific data from a pre-trained model while preserving the knowledge derived from other data [28]. The most straightforward approach, known as exact unlearning, involves retraining the model from scratch using a refined dataset [26] (hereafter referred to as the “retrained model”). While this approach is regarded as the gold standard because it guarantees complete elimination of any influence from data designated for removal, it incurs significant computational expenses due to the necessity of retraining models on refined datasets whenever unlearning is required. To overcome these limitations, several approximate unlearning algorithms have been proposed, offering more efficient alternatives while striving to maintain the goal of machine unlearning [16]. In image classification tasks, approximate unlearning algorithms are typically evaluated based on their ability to produce models that achieve classification accuracy similar to that of the exact unlearning on the given dataset (e.g., retain and forget sets) [28], as illustrated on the left side of Fig. 1. Following this evaluation method, several methods have induced unlearning through techniques such as using Random Labeling [13], Gradient Ascent, [32] and metric learning [2], while mitigating the over-forgetting by incorporating additional regularization terms [23, 12, 5, 2].
@@
Figure 1: A comparison of (a) the traditional evaluation framework and (b) our proposed evaluation framework. Note that $\theta_{u}$ and $\theta_{r}$ refer to the unlearned model and the model trained on the retain set, respectively. Traditional unlearning evaluation methods primarily focus on analyzing the unlearned model’s output logits to assess the effectiveness of unlearning under small-scale scenarios such as CIFAR-10. In contrast, our framework introduces additional evaluation factors by examining the unlearned model’s feature representation similarity with $\theta_{r}$ in terms of transferability and representational similarity under the large-scale unlearning scenario, such as using ImageNet-1K.
@@
Figure 2: Performance comparison between logit-based (left) and representation-based evaluation (right) reveals contrasting findings.
@@
In this paper, we propose a comprehensive new evaluation framework that addresses the limitations of existing small-scale, logit-based evaluations by incorporating assessments of unlearned models ($\theta_{u}$) from the perspective of feature representations in large-scale settings. As illustrated in Fig. 1(b), our framework evaluates the representations of unlearned models using Centered Kernel Alignment (CKA) [20] to compare the feature representations of the unlearned models with those of a model trained only on the retain set ($\theta_{r}$). Additionally, we apply $k$-Nearest Neighbors ($k$-NN) analysis on the feature space across various downstream datasets to assess the quality of the unlearned model’s feature representations and compare the results with those of $\theta_{r}$. Through these evaluation methods, we re-evaluate multiple state-of-the-art unlearning algorithms in the large-scale unlearning scenario. Fig. 2 illustrates key findings from experiments on unlearning 100 random classes in ImageNet-1K using ResNet-50: while logit-based evaluation (e.g., classification accuracy) demonstrates successful unlearning by showing a decreasing gap between the unlearned models ($\theta_{u}$) and the retrained model ($\theta_{r}$) (Fig. 2, left), a representation-based evaluation using CKA reveals a critical limitation (Fig. 2, right). In an ideal scenario, where unlearning is truly effective, the unlearned model should be indistinguishable from the retrained model, with the ideal value being 10. However, as shown in Fig. 2 (right), all unlearned models exhibit greater similarity to the original model than to the retrained model. This highlights a fundamental issue: existing unlearning methods fail to sufficiently erase learned information at the representation level. This reinforces the need for representation-based evaluation to rigorously assess whether unlearning truly eliminates the influence of the targeted data.
Furthermore, random class forgetting has been widely adopted as a standard evaluation scenario in unlearning studies [5, 12, 2, 23]. However, we highlight a critical limitation of this approach: the forgetting process may be inadequately assessed, as similar representational knowledge can persist across multiple alternative classes, making it difficult to verify whether the targeted data has been truly unlearned. To address this issue, we propose a novel unlearning evaluation paradigm, Top Class-wise Forgetting, in which the classes selected for unlearning are semantically similar to those in downstream tasks. This setup requires a substantial divergence of feature representations from the original model, offering a more rigorous assessment of the unlearning algorithm’s effectiveness. We then evaluate existing unlearning methods using both logit-based and representation-based metrics across diverse downstream datasets. Our goal is to establish a reusable and scalable benchmark for the unlearning community, enabling standardized and holistic evaluation of future methods. In summary, the core contributions are as follows:
  • We identify fundamental limitations in current unlearning evaluations, which overly rely on logit-based metrics under small-scale settings.
  • To address this, we propose a unified benchmark framework that incorporates representation-based metrics and reveals that many unlearning methods fail to alter internal representations.
  • We further introduce Top Class-wise Forgetting, a more rigorous evaluation scenario that emphasizes semantic overlap with downstream tasks and better reflects real-world unlearning challenges.
@@ 2 Related Works
Machine unlearning in image classification. Early research on machine unlearning in neural networks has primarily focused on image classification, leading to the development of various unlearning algorithms with distinct strategies [28]. A widely used approach, Gradient Ascent (GA), maximizes the loss on the forget set to erase its influence, making it one of the most fundamental yet widely adopted techniques in recent studies [14, 31, 15, 37, 5]. Similarly, Fine-tuning achieves unlearning by inducing catastrophic forgetting on forget set, updating the model solely on retain set [14, 36]. More sophisticated unlearning techniques have been developed: SCRUB [23] employs a knowledge distillation framework to selectively forget knowledge from the forget set while retaining other knowledge. L2UL [5] enhances unlearning by incorporating two regularization terms, which mitigate forgetting at both the feature and weight levels when combined with GA. SalUn [12] modifies the most influential parameters through weight saliency, enabling selective unlearning. SCAR [2] leverages metric learning alongside a distillation-based strategy, aligning forget set representations to the nearest incorrect class using the Mahalanobis distance.
@@
Evaluation and analysis of representations. Transfer learning has become a foundational technique in machine learning, facilitating the adaptation of pre-trained models to diverse downstream tasks [29]. Beyond its practical applications, it is widely used to assess the quality of the learned representations. For instance, it has been employed to compare the learned representations of different model architectures [21] and algorithms across various domains, such as self-supervised learning [8, 17] and continual learning [6]. This evaluation is typically performed using metrics such as $k$-Nearest Neighbors ($k$-NN), which measure how well the learned representations generalize in various downstream tasks. Centered Kernel Alignment (CKA) [20] has been widely used to compare feature representations across models, providing insights into their structural similarities and differences. Furthermore, t-SNE [33] has been extensively utilized to visualize feature representations, offering an intuitive means to compare and analyze how different models organize learned representations.
In contrast to prior work, we revisit the foundations of unlearning evaluation by exposing key limitations in conventional protocols and propose a unified, representation-aware benchmark that enables more realistic and rigorous assessment.
3 Why We Need Better Benchmarks: An Empirical Study
@@
In this section, we present a comprehensive analysis of existing machine unlearning methods, focusing on both logit-based and representation-based evaluations in the large-scale setup. We review the classical unlearning evaluation protocols and introduce a new evaluation metric and setup in Sec. 3.1. Then, we introduce the motivation of our approach by highlighting limitations in current evaluation frameworks in Sec. 3.2, introduce our proposed evaluation metric in Sec. 4, present experimental findings under large-scale settings from a representation perspective in Sec. 4.1, and show results under top class-wise forgetting scenarios from a transfer learning perspective in Sec. 4.2.
3.1 Problem Setup and Evaluation Metrics for Unlearning
@@
Problem setup and objective of machine unlearning in the classification task.   We define the entire dataset as $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, where $x_{i}$ represents the input image, $y_{i}$ denotes its corresponding label, and $N$ is the total number of samples in the dataset. The subset of data targeted for unlearning, referred to as the forget set, is denoted by $\mathcal{D}_{f}\subset\mathcal{D}$, and consists of samples from specific classes that the model is intended to forget. Retain set, denoted by $\mathcal{D}_{r}=\mathcal{D}\setminus\mathcal{D}_{f}$, includes all samples in $\mathcal{D}$ that are not part of the forget set. Additionally, we introduce the test forget set, denoted as $\mathcal{D}^{\text{te}}_{f}$, which consists of data points from the same classes as $\mathcal{D}_{f}$ but not included in the train set. Similarly, the test retain set, denoted as $\mathcal{D}^{\text{te}}_{r}$, contains data points from the same classes as $\mathcal{D}_{r}$. We assume that the original model $\theta$ is trained on the entire dataset $\mathcal{D}$ and subsequently unlearns information related to the forget set $\mathcal{D}_{f}$. Note that the goal of machine unlearning is to produce an unlearned model $\theta_{u}$ that retains the knowledge of $\mathcal{D}_{r}$ while achieving the unlearning of $\mathcal{D}_{f}$. Additionally, we consider a retrained model $\theta_{r}$, which is trained on $\mathcal{D}_{r}$ from scratch, as an oracle for comparison in the unlearning process.
@@
Logit-based evaluation metrics.    As a classical evaluation protocol, unlearning models are evaluated on the following metrics: Forget Accuracy (FA): Measures the accuracy of the unlearned model $\theta_{u}$ on $\mathcal{D}_{f}$. Retain Accuracy (RA): Measures the accuracy of the unlearned model $\theta_{u}$ on $\mathcal{D}_{r}$. Test Forget Accuracy (TFA): Classification accuracy on the test forget set $\mathcal{D}^{\text{te}}_{f}$. Similar to FA, this evaluates the model’s ability to forget removed classes beyond the training data. Test Retain Accuracy (TRA): Classification accuracy on the test retain set $\mathcal{D}^{\text{te}}_{r}$. It indicates how effectively the model’s knowledge generalizes to unseen data of the remaining classes. Following [39], we introduce the Average GAP of Logit-based Metrics (AGL), which quantifies the discrepancy between the given model (e.g., the unlearned model $\theta_{u}$) and the golden standard $\theta_{r}$. This metric aggregates the logit-based evaluations:
$AGL=$ $\Bigl{(}1-\mathrm{G}(\theta_{u},\mathcal{D}_{f})\Bigr{)}\times\Bigl{(}1-\mathrm{G}(\theta_{u},\mathcal{D}_{r})\Bigr{)}\times\Bigl{(}1-\mathrm{G}(\theta_{u},\mathcal{D}^{\text{te}}_{f})\Bigr{)}\times\Bigl{(}1-\mathrm{G}(\theta_{u},\mathcal{D}^{\text{te}}_{r})\Bigr{)},$ (1)
@@
where $\mathcal{D}_{f}$, $\mathcal{D}_{r}$, $\mathcal{D}^{\text{te}}_{f}$, and $\mathcal{D}^{\text{te}}_{r}$ represent the source training datasets and their corresponding test sets, and $\mathrm{G}$ computes the classification accuracy differences, $\mathrm{G}(\theta_{u},\mathcal{D}_{\text{}})=\bigl{|}\,\text{A}(\theta_{u},\mathcal{D}_{\text{}})-\text{A}(\theta_{r},\mathcal{D}_{\text{}})\bigr{|}$, where $\mathrm{A}$ represents the classification accuracy on $\mathcal{D}$.
@@
Figure 3: We visualize the feature representations from the original ($\theta_{o}$) retrained ($\theta_{r}$), and unlearned models ($\theta_{u}$) on a subset of ImageNet-1K using ResNet-50. Unlike $\theta_{r}$, which serves as the gold standard, the forget classes (highlighted in red) in the existing unlearning baselines (e.g., (d), (e), and (f)) are severely dispersed throughout the entire class distribution.
Representation-based evaluation metrics.   To overcome the limitations of logit-based evaluation, we propose an additional representation-based approach that assesses unlearning effectiveness by comparing features extracted from the encoder (e.g., after the average pooling layer in ResNet) between the unlearned and retrained models. This enables a fine-grained analysis of representational changes beyond what output-level metrics can capture. We use the following two metrics for the representation-based evaluation: First, we use Centered Kernel Alignment (CKA) [20] to measure the similarity between the feature representations of $\theta_{u}$ and $\theta_{r}$ as well as $\theta_{u}$ and $\theta_{o}$. Second, we measure $k$-Nearest Neighbors Accuracy ($k$-NN) [9] to assess the representational quality of $\theta_{u}$. We employ a $k$-NN classifier, which takes features of the test data from a model and performs classification based on the neighborhood features. Additionally, we also employ t-SNE [33] to visualize feature distributions. More details are provided in Sec. A.2 in the Appendix.
@@
Unlearning baselines.   We evaluate several representative machine unlearning algorithms in the classification task, which can be categorized into two groups: 1) Unlearning without $\mathcal{D}_{r}$, including Gradient Ascent (GA)[32], Random Labeling (RL)[13], Pseudo Labeling (PL)[7], and SalUn[12]; and 2) unlearning with $\mathcal{D}_{r}$, including DUCK [11], Contrastive Unlearning (CU) [38], SCRUB [23], and SCAR [2]. Note that SalUn and SCAR are known as current state-of-the-art methods in each category. A more detailed explanation of each baseline can be found in Sec. A.1 in the Appendix.
Unlearning scenarios.   In our experiments and analysis, we consider two unlearning scenarios. The first is Random-$N$ Class-wise Forgetting, where we randomly select $N$ classes to remove. This scenario is a standard approach to evaluate unlearning algorithms. To provide a more comprehensive assessment, we introduce a second scenario, Top-$N$ Class-wise Forgetting, in which we specifically target classes that exhibit semantic similarity to the downstream task classes. The selection of $N$ classes is based on their relevance to the downstream task.